OpenStudy (anonymous):

(z^2)/(2z+14) multiply (2z+14)/(2z^6)

OpenStudy (anonymous):

\[\frac{z^2}{2z + 14} \times \frac{2z + 14}{2z^6}\] note (2z + 14) cancels, and z^2 will cancel partially with 2z^6, leaving: \[\frac{1}{2z^4} \]

OpenStudy (anonymous):

thanks again I'm a bit hopeless at this, right now anyway

OpenStudy (anonymous):

Don't give up! Let me know what I wrote makes sense, or you need more help

OpenStudy (anonymous):

yes I understand it very well